# Latent Factor Analysis in Short Panels

### v2

@@ 4.3 Spherical errors
If errors are spherical, i.e., matrix $V_{\varepsilon}=\bar{\sigma}^{2}I_{T}$ is a multiple of the identity with unknown parameter $\bar{\sigma}^{2}>0$, then the asymptotic distribution of $LR(k)$ corresponds to a special case of Proposition 2 (a). If sphericity is imposed in the estimation procedure, i.e., $\hat{V}_{\varepsilon}$ becomes $\hat{V}_{\varepsilon,c}=\hat{\sigma}^{2}I_{T}$, the constrained FA estimator boils down to the Principal Component Analysis (PCA) estimator; see Anderson and Rubin (1956) Section 7.3. Let $\hat{F}_{c}$ denote the FA estimator under sphericity constraint. Then, $\hat{F}_{c}$ is the matrix of eigenvectors of matrix $\hat{V}_{y}$ standardized such that $\hat{F}_{c}^{\prime}\hat{F}_{c}=diag(\hat{\delta}_{1}-\hat{\sigma}^{2},...,\hat{\delta}_{k}-\hat{\sigma}^{2})$, and $\hat{\sigma}^{2}=\frac{1}{T-k}\sum_{j=k+1}^{T}\hat{\delta}_{j}$, where $\hat{\delta}_{j}$ denotes the $j$th largest eigenvalue of matrix $\hat{V}_{y}$. The matrix $\hat{S}$ under sphericity constraint becomes $\hat{S}_{c}=\frac{1}{\hat{\sigma}^{2}}M_{\hat{F}_{c}}(\hat{V}_{y}-\hat{\sigma}^{2}I_{T})M_{\hat{F}_{c}}=\frac{1}{\hat{\sigma}^{2}}\left(\frac{1}{n}\hat{\varepsilon}_{c}\hat{\varepsilon}_{c}^{\prime}\right)-M_{\hat{F}_{c}},$ where $M_{\hat{F}_{c}}=I_{T}-\hat{F}_{c}(\hat{F}_{c}^{\prime}\hat{F}_{c})^{-1}\hat{F}^{\prime}$ and $\hat{\varepsilon}_{c}=M_{\hat{F}_{c}}\tilde{Y}$ is the matrix of OLS residuals. Then, the constrained LR statistic becomes $LR_{c}(k):=-n\sum_{j=k+1}^{T}\log(1+\frac{\hat{\delta}_{j}-\hat{\sigma}^{2}}{\hat{\sigma}^{2}})$ and reduces to the LR statistic invoqued by Onatski (2023) in his discussion of FGS. [^18] We have $\hat{\gamma}_{j}=\hat{\delta}_{j}/\hat{\sigma}^{2}-1$, and $LR_{c}(k)=\frac{n}{2\hat{\sigma}^{4}}\sum_{j=k+1}^{T}(\hat{\delta}_{j}^{2}-\hat{\sigma}^{2})^{2}+o_{p}(1)$, i.e., the constrained LR statistic is asymptotically equivalent to the sum of squared deviations of the $T-k$ smallest eigenvalues from their mean. Besides, by similar results, we have that the eigenvalue spacing statistic $\sqrt{n}\mathscr{S}(k):=\hat{\gamma}_{k+1}-\hat{\gamma}_{T}$ corresponds to the statistic considered in FGS divided by $\hat{\sigma}^{2}$, and its asymptotic distribution under sphericity coincides with that obtained by FGS. Under spherical errors, Onatski (2023) shows $LR_{c}(k)\Rightarrow\frac{1}{2}(Tr[(Z^{*})^{2}]-\frac{1}{T-k}[Tr(Z^{*})]^{2})=vech(Z^{*})^{\prime}M_{\boldsymbol{x}}vech(Z^{*})$, where $M_{\boldsymbol{x}}=I_{p}-\boldsymbol{x}(\boldsymbol{x}^{\prime}\boldsymbol{x})^{-1}\boldsymbol{x}^{\prime}$ with $\boldsymbol{x}=vech(I_{T-k})$, $Z^{*}=\frac{1}{\bar{\sigma}^{4}}G^{\prime}ZG$, $Z$ is the distributional limit of $Z_{n}$, and $G$ is a $T\times(T-k)$ matrix such that $F^{\prime}G=0$ and $G^{\prime}G=\bar{\sigma}^{2}I_{T-k}$. Hence the asymptotic distribution of the constrained LR statistic under sphericity is $\sum_{j=1}^{p-1}\mu_{j}\chi^{2}(1)$, where the $\chi^{2}(1)$ are independent and the $\mu_{j}$ are the non-zero eigenvalues of matrix $M_{\boldsymbol{x}}\Omega_{Z^{*}}M_{\boldsymbol{x}}$, with $\Omega_{Z^{*}}=V[vech(Z^{*})]$. Under Gaussian errors, it simplifies to $LR_{c}(k)\Rightarrow q\chi^{2}(p-1)$ (see Section 4.1). In a PCA setting when sphericity fails to hold, a massive over-rejection by over 80 percentage points for a nominal size of 5% is again observed in our Monte Carlo experiments. It is true both for the constrained LR test and for the eigenvalue spacing test of FGS. We get an average estimated number of factors often above 10 instead of 2 when $T=24$.
@@ 5 Local asymptotic power
In this section, we study the asymptotic power of the test statistics against local alternatives in which we have $k$ (strong) factors plus a weak factor. Specifically, under $H_{1,loc}(k)$, we have $\sqrt{n}\gamma_{k+1}\rightarrow c_{k+1}$ as $n\rightarrow\infty$, with $c_{k+1}>0$. The (drifting) DGP is $Y=\mu 1_{n}^{\prime}+F\beta^{\prime}+F_{k+1}\beta_{loc}^{\prime}+\varepsilon$, where $\beta_{loc}$ is the loading vector for the $(k+1)$th factor, and the factor vector is normalized such that $F_{k+1}=\sqrt{\gamma_{k+1}}\rho_{k+1}$ with $\rho_{k+1}^{\prime}V_{\varepsilon}^{-1}\rho_{k+1}=1$ and $F^{\prime}V_{\varepsilon}^{-1}\rho_{k+1}=0$. Thus, we can write $\rho_{k+1}=G\xi_{k+1}$ for a $T-k$ dimensional vector $\xi_{k+1}$ with unit norm. Scalar $c_{k+1}$ and vector $\xi_{k+1}$ yield the (normalized) strength and the direction of the local alternative.
@@ 5.1 Asymptotic distributions under local alternatives
We derive an asymptotic expansion of $n\|\hat{S}\|^{2}$ under $H_{1,loc}(k)$ using similar arguments as in the proof of Proposition 2 (a) (see the proof of Proposition 4 in Appendix B for the derivation):
$\frac{n}{2}\|\hat{S}\|^{2}=vech(Z_{n,loc}^{*})^{\prime}M_{\boldsymbol{X}}vech(Z_{n,loc}^{*})+o_{p}(1),$ (6)
where $Z_{n,loc}^{*}=Z^{*}_{n}+c_{k+1}\xi_{k+1}\xi_{k+1}^{\prime}$. From the CLT, we have $Z_{n,loc}^{*}\Rightarrow Z_{loc}^{*}$ where $Z_{loc}^{*}=Z^{*}+c_{k+1}\xi_{k+1}\xi_{k+1}^{\prime}$. Matrix variate $Z^{*}_{loc}$ is a non-central symmetric Gaussian matrix. The non-zero mean depends in general on both $c_{k+1}$ and $\xi_{k+1}$, while the variances and covariances of the elements of $Z^{*}_{loc}$ are the same as those of $Z^{*}$. Then, we deduce from (6) that the asymptotic distribution of the $LR(k)$ statistic under the local alternative hypothesis is a weighted average of $df$ mutually independent non-central chi-square variables, with non-centrality parameters depending on $vech(\Delta):=M_{\boldsymbol{X}}vech(c_{k+1}\xi_{k+1}\xi^{\prime}_{k+1})$.
@@
Let Assumptions 1-2, A.1-A.6 hold. Under the local alternative hypothesis $H_{1,loc}(k)$, we have as $n\rightarrow\infty$ and $T$ is fixed, $LR(k)\Rightarrow\sum_{j=1}^{df}\mu_{j}\chi^{2}(1,\lambda_{j}^{2})$, where $\lambda_{j}^{2}=\mu_{j}^{-1}[v_{j}^{\prime}vech(\Delta)]^{2}$, and the $\mu_{j}$ and $v_{j}$ are the non-zero eigenvalues and the associated standardized eigenvectors of matrix $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}$.
The non-centrality term $vech(\Delta)$ is in charge of the asymptotic local power of the statistics. When this vector is null, the asymptotic local power is zero. Indeed, for some local alternatives the $(k+1)$th weak factor can be absorbed in the diagonal variance matrix $V_{\varepsilon}$ of the error terms. More precisely, in Appendix E.3 ii), we show that $V_{y}+\frac{c_{k+1}}{\sqrt{n}}\rho_{k+1}\rho_{k+1}^{\prime}=F^{*}(F^{*})^{\prime}+V_{\varepsilon}^{*}+\frac{1}{\sqrt{n}}G\Delta G^{\prime}+o(1/\sqrt{n})$ for some $T\times k$ matrix $F^{*}$ and diagonal matrix $V_{\varepsilon}^{*}$, which yields asymptotically a $k$-factor model when $\Delta=0$. We have $\lambda^{2}:=\sum_{j=1}^{df}\mu_{j}\lambda_{j}^{2}=\frac{1}{2}\|\Delta\|^{2}$, i.e., the half squared Frobenius norm of the matrix measuring local distance from the $k$-factor specification. It follows that the asymptotic local power of the LR statistic is non null as long as $\lambda^{2}>0$, i.e., it has non-trivial asymptotic power against any proper local alternative. In our Monte Carlo experiments in Appendix C, we find that the LR statistic has size close to the nominal value, and power against global as well as local alternatives with time dimension as small as $T=6$.
Under the normality of errors, or more generally the conditions of Proposition 3, using that matrix $\frac{1}{\sqrt{\bar{q}}}Z^{*}$ is in the GOE for dimension $T-k$, i.e., $vech(Z^{*})\sim N(0,\bar{q}I_{p})$, we have $LR(k)\Rightarrow\bar{q}\chi^{2}(df,\lambda^{2}/\bar{q})$ from (6). The local power is a function solely of the squared Euclidean norm of the vector $vech(\Delta)$ measuring local distance from the $k$-factor specification, divided by $\bar{q}$.
@@ 5.2 AUMPI tests
In this subsection, we investigate asymptotic local optimality of the LR statistic for testing hypotheses on the number of latent factors. In our framework with composite null and alternative hypotheses and multi-dimensional parameter, we cannot expect in general to establish Uniformly Most Powerful (UMP) tests. Instead, we can establish an optimality property by restricting the class of tests to invariant tests (e.g. Lehmann and Romano (2005)). We focus on statistics with test functions $\phi$ written on the elements of matrix $\hat{S}$. To eliminate the asymptotic redundancy in the elements of $\hat{S}$, we actually consider the test class $\mathscr{C}=\left\{\phi~{}:~{}\phi=\phi(\hat{W})\right\}$ with $\hat{W}:=\sqrt{n}\hat{\boldsymbol{D}}^{\prime}vech(\hat{S}^{*})$, where $\hat{S}^{*}=\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1/2}\hat{S}\hat{V}_{\varepsilon}^{-1/2}\hat{G}$ and $\hat{\boldsymbol{D}}$ is a $p\times df$ full-rank matrix, such that $M_{\hat{\boldsymbol{X}}}=\hat{\boldsymbol{D}}\hat{\boldsymbol{D}}^{\prime}$ and $\hat{\boldsymbol{D}}^{\prime}\hat{\boldsymbol{D}}=I_{df}$. We have that $\hat{S}^{*}=diag(\hat{\gamma}_{k+1},\dots,\hat{\gamma}_{T})$, since $\hat{S}$ has the eigendecomposition $\hat{V}_{\varepsilon}^{-1/2}\hat{G}diag(\hat{\gamma}_{k+1},\dots,\hat{\gamma}_{T})\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1/2}$ from Proposition 1 (a). The diagonal matrix $\hat{S}^{*}$ contains the information in $\hat{S}$ beyond orthogonality of its rows and columns to $\hat{V}_{\varepsilon}^{-1/2}\hat{F}$. The vector $\hat{W}$ contains the information in $\sqrt{n}vech(\hat{S}^{*})$ beyond orthogonality to $\hat{\boldsymbol{X}}$. [^19] From Proposition 1 (c), we have $0=diag(\hat{S})=2\hat{V}^{-1}_{\varepsilon}\hat{\boldsymbol{X}}^{\prime}vech(\hat{S}^{*})$. Therefore, $vech(\hat{S}^{*})$ lies in the orthogonal complement of the range of $\hat{\boldsymbol{X}}$ in sample.
@@
Matrices $\hat{G}$ and $\hat{\boldsymbol{D}}$ are both consistent up to post-multiplication by an orthogonal matrix. This point yields a group of orthogonal transformations under which we require the test statistics to be invariant. [^20] Here, we do not deal with invariance to data transformations but rather with invariance to parameterization of $\hat{G}$ and $\hat{\boldsymbol{D}}$. However, if we consider tests based on the elements of vector $\hat{W}$, this difference is immaterial. In Appendix E.6, we show that the maximal invariant under this group is provided by $\hat{W}^{\prime}\hat{W}=nvech(\hat{S}^{*})^{\prime}M_{\hat{\boldsymbol{X}}}vech(\hat{S}^{*})$. Since $\sqrt{n}vech(\hat{S}^{*})$ belongs to the range of matrix $M_{\hat{\boldsymbol{X}}}$, we have $\hat{W}^{\prime}\hat{W}=\frac{n}{2}\|\hat{S}^{*}\|^{2}=\frac{n}{2}\|\hat{S}\|^{2}$. Therefore, the invariant tests are functions of the squared norm of $\hat{S}$, which is asymptotically equivalent to the LR statistic (up to the factor $1/2$).
@@
In the Gaussian case, or more generally under the conditions of Proposition 3, the LR statistic follows asymptotically a scaled non-central chi-square distribution with $df$ degrees of freedom and non-centrality parameter $\lambda^{2}/\bar{q}=\sum_{j=1}^{df}\lambda_{j}^{2}$ as shown in the previous subsection. Thus, we can simplify the null and alternative hypotheses of our testing problem asymptotically and locally to a one-sided test with null hypothesis $H_{0}(k):\lambda^{2}=0$ vs. alternative hypothesis $H_{1,loc}(k):\lambda^{2}>0$. The scaling constant $q>0$ plays no role in the power analysis. It means that the LR test is an AUMPI test (Lehmann and Romano (2005) Chapters 3 and 13). Indeed, the density $g(z;df,\lambda^{2})$ of the $\chi^{2}(df,\lambda^{2})$ distribution is Totally Positive of order 2 ($TP2$) in $z$ and $\lambda^{2}$ (Eaton (1987) Example A.1 p. 468); see Miravete (2011) for a review of applications of TP2 in economics. A density, which is $TP2$ in $z$ and $\lambda^{2}$, has the Monotone Likelihood Ratio (MLR) property (Eaton (1987) p. 467). Since $g(z;df,\lambda^{2})/g(z;df,0)$ is an increasing function in $z$, it gives the AUMPI property.
@@
In the general case with $df>1$, when neither Gaussianity nor the conditions of Proposition 3 apply, we cannot use the same reasoning, since the density $f(z;\lambda_{1},...,\lambda_{df})$ of $\sum_{j=1}^{df}\mu_{j}\chi^{2}(1,\lambda_{j}^{2})$, with $\mu_{j}>0$, $j=1,...,df$, is not a function of $\lambda^{2}=\sum_{j=1}^{df}\mu_{j}\lambda_{j}^{2}$ only, and thus cannot be $TP2$ in $z$ and $\lambda^{2}$. Instead, our strategy to get a new AUMPI result uses a power series representation of the density of $\sum_{j=1}^{df}\mu_{j}\chi^{2}(1,\lambda_{j}^{2})$ in terms of central chi-square densities from Kotz, Johnson, and Boyd (1967). Under the sufficient condition (7) in Proposition 5, the density ratio $\frac{f(z;\lambda_{1},...,\lambda_{df})}{f(z;0,...,0)}$ is monotone increasing in $z$.
@@
Let Assumptions 1-2, A.1-A.6 hold. (a) Let us assume that, for any DGP in the subset $\bar{H}_{1,loc}(k)\subset H_{1,loc}(k)$ of the local alternative hypothesis, we have for any integer $m\geq 3$:

### v3

@@ 5.3 Spherical errors
If errors are spherical, i.e., matrix $V_{\varepsilon}=\bar{\sigma}^{2}I_{T}$ is a multiple of the identity with unknown parameter $\bar{\sigma}^{2}>0$, then the asymptotic distribution of $LR(k)$ corresponds to a special case of Proposition 1 (a). If sphericity is imposed in the estimation procedure, i.e., $\hat{V}_{\varepsilon}$ becomes $\hat{V}_{\varepsilon,c}=\hat{\sigma}^{2}I_{T}$, the constrained FA estimator boils down to the Principal Component Analysis (PCA) estimator; see Anderson and Rubin (1956) Section 7.3. If we have independent individual ARCH processes across assets and no common ARCH effects, PCA is still valid since one inherits sphericity from cross-sectional averaging. Let $\hat{F}_{c}$ denote the FA estimator under sphericity constraint. Then, $\hat{F}_{c}$ is the matrix of eigenvectors of matrix $\hat{V}_{y}$ standardized such that $\hat{F}_{c}^{\prime}\hat{F}_{c}=diag(\hat{\delta}_{1}-\hat{\sigma}^{2},...,\hat{\delta}_{k}-\hat{\sigma}^{2})$, and $\hat{\sigma}^{2}=\frac{1}{T-k}\sum_{j=k+1}^{T}\hat{\delta}_{j}$, where $\hat{\delta}_{j}$ denotes the $j$th largest eigenvalue of matrix $\hat{V}_{y}$. Then, the constrained LR statistic becomes $LR_{c}(k):=-n\sum_{j=k+1}^{T}\log(1+\frac{\hat{\delta}_{j}-\hat{\sigma}^{2}}{\hat{\sigma}^{2}})$ and reduces to the LR statistic invoqued by Onatski (2023) in his discussion of FGS. We have $\hat{\gamma}_{j}=\hat{\delta}_{j}/\hat{\sigma}^{2}-1$, and $LR_{c}(k)=\frac{n}{2\hat{\sigma}^{4}}\sum_{j=k+1}^{T}(\hat{\delta}_{j}^{2}-\hat{\sigma}^{2})^{2}+o_{p}(1)$, i.e., the constrained LR statistic is asymptotically equivalent to the sum of squared deviations of the TkT-k smallest eigenvalues from their mean. Under spherical errors, Onatski (2023) shows $LR_{c}(k)\Rightarrow\frac{1}{2}(Tr[(Z^{*})^{2}]-\frac{1}{T-k}[Tr(Z^{*})]^{2})=vech(Z^{*})^{\prime}M_{\boldsymbol{x}}vech(Z^{*})$, where $M_{\boldsymbol{x}}=I_{p}-\boldsymbol{x}(\boldsymbol{x}^{\prime}\boldsymbol{x})^{-1}\boldsymbol{x}^{\prime}$ with $\boldsymbol{x}=vech(I_{T-k})$, $^{*}=\frac{1}{\bar{\sigma}^{4}}G^{\prime}ZG, ZZ$ is the distributional limit of $Z_{n}$, and $G$ is a $T\times(T-k)$ matrix such that $F^{\prime}G=0$ and $G^{\prime}G=\bar{\sigma}^{2}I_{T-k}$. Hence the asymptotic distribution of the constrained LR statistic under sphericity is driven by $\sum_{j=1}^{p-1}\mu_{j}\chi^{2}_{j}(1)$, where the $\chi^{2}_{j}(1)$ are independent and the $\mu_{j}$ are the non-zero eigenvalues of matrix $M_{\boldsymbol{x}}\Omega_{Z^{*}}M_{\boldsymbol{x}}$, with $\Omega_{Z^{*}}=V[vech(Z^{*})]$. Under Gaussian errors, it simplifies to $LR_{c}(k)\Rightarrow q\chi^{2}(p-1)$ (see Section 5.1). In a PCA setting when sphericity fails to hold, a massive over-rejection by over 80 percentage points for a nominal size of 5% is again observed in our Monte Carlo experiments. It is true both for the constrained LR test and for the eigenvalue spacing test of FGS. We get an average estimated number of factors often above 10 instead of 2 when $T=24$.
@@ 6 Local asymptotic power
In this section, we study the asymptotic power properties of the test statistics against local alternatives in which we have $k$ (strong) factors plus a weak factor. [^1] Recently, Barigozzi and Hallin (2025) show that exchangeability of assets in the whole cross-section rules out weak factors as defined in this section. In equity markets, ordering matters because of industry sectors and other firm characteristics such as size and book-to-market. Weak factors can result from strong factors that load only on a subsample of assets of size $O(\sqrt{n})$ or from factors with vanishing loadings of size $O(n^{-1/4})$ diluted in the cross-section. Specifically, under $H_{1,loc}(k)$, we have $\sqrt{n}\gamma_{k+1}\rightarrow c_{k+1}$ as $n\rightarrow\infty$, with $c_{k+1}>0$. The (drifting) DGP is $Y=\mu 1_{n}^{\prime}+F\beta^{\prime}+F_{k+1}\beta_{loc}^{\prime}+\varepsilon$, where $\beta_{loc}$ is the loading vector for the $(k+1)$th factor, and the factor vector is normalized such that $F_{k+1}=\sqrt{\gamma_{k+1}}\rho_{k+1}$ with $\rho_{k+1}^{\prime}V_{\varepsilon}^{-1}\rho_{k+1}=1$ and $F^{\prime}V_{\varepsilon}^{-1}\rho_{k+1}=0$. Thus, we can write $\rho_{k+1}=G\xi_{k+1}$ for a TkT-k dimensional vector $\xi_{k+1}$ with unit norm. Scalar $c_{k+1}$ and vector $\xi_{k+1}$ yield the (normalized) strength and the direction of the local alternative. The root-$n$ condition defining $H_{1,loc}(k)$ represents the weakest level of factor strength that our test can detect. This detection threshold matches the convergence rate of the FA estimators. Intuitively, vanishing factors satisfying $\sqrt{n}\gamma_{k+1}\to 0$ cannot be detected because their signal is dominated by estimation noise. In contrast, the test achieves asymptotic power of 1 against semi-strong factors for which $\sqrt{n}\gamma_{k+1}\to\infty$, because their signal dominates estimation noise. The hedge case $H_{1,loc}(k)$ corresponds to the signal being of the same order as the estimation noise, for which asymptotic power is between 0 and 1.
@@ 6.1 Asymptotic distributions under local alternatives
We derive an asymptotic expansion of the LR statistics under $H_{1,loc}(k)$ using similar arguments as in the proof of Proposition 1 (a) (see the proof of Proposition 3 in Appendix A for the derivation):
$LR(k)=vech(Z_{n,loc}^{*})^{\prime}M_{\boldsymbol{X}}vech(Z_{n,loc}^{*})+o_{p}(1),$ (4)
where $Z_{n,loc}^{*}=Z^{*}_{n}+c_{k+1}\xi_{k+1}\xi_{k+1}^{\prime}$. From the CLT, we have $Z_{n,loc}^{*}\Rightarrow Z_{loc}^{*}$ where $Z_{loc}^{*}=Z^{*}+c_{k+1}\xi_{k+1}\xi_{k+1}^{\prime}$. Matrix variate $Z^{*}_{loc}$ is a non-central symmetric Gaussian matrix. The non-zero mean depends in general on both $c_{k+1}$ and $\xi_{k+1}$, while the variances and covariances of the elements of $Z^{*}_{loc}$ are the same as those of $Z^{*}$. Then, we deduce from (4) that the asymptotic distribution of the $LR(k)$ statistic under the local alternative hypothesis is a weighted sum of $df$ mutually independent non-central chi-square variables, with non-centrality parameters depending on $vech(\Delta):=M_{\boldsymbol{X}}vech(c_{k+1}\xi_{k+1}\xi^{\prime}_{k+1})$.
@@
Let Assumptions 1-8 hold. Under the local alternative hypothesis $H_{1,loc}(k)$, we have as $n\rightarrow\infty$ and $T$ is fixed, $LR(k)\Rightarrow\sum_{j=1}^{df}\mu_{j}\chi^{2}(1,\lambda_{j}^{2})$, where $\lambda_{j}^{2}=\mu_{j}^{-1}[v_{j}^{\prime}vech(\Delta)]^{2}$, and the $\mu_{j}$ and $v_{j}$ are the non-zero eigenvalues and the associated standardized eigenvectors of matrix $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}$.
The non-centrality term $vech(\Delta)$ is in charge of the asymptotic local power of the statistics. When this vector is null, the asymptotic local power is zero. Indeed, for some local alternatives the $(k+1)$th weak factor can be absorbed in the diagonal covariance matrix $V_{\varepsilon}$ of the error terms. More precisely, in Appendix B.4 ii) we show that $V_{y}+\frac{c_{k+1}}{\sqrt{n}}\rho_{k+1}\rho_{k+1}^{\prime}=F^{*}(F^{*})^{\prime}+V_{\varepsilon}^{*}+\frac{1}{\sqrt{n}}G\Delta G^{\prime}+o(1/\sqrt{n})$ for some $T\times k$ matrix $F^{*}$ and diagonal matrix $V_{\varepsilon}^{*}$, which yields asymptotically a $k$-factor model when $\Delta=0$. We have $\lambda^{2}:=\sum_{j=1}^{df}\mu_{j}\lambda_{j}^{2}=\frac{1}{2}\|\Delta\|^{2}$, i.e., the half squared Frobenius norm of the matrix measuring local distance from the $k$-factor specification. It follows that the asymptotic local power of the LR statistic is non null as long as $\lambda^{2}>0$, i.e., it has non-trivial asymptotic power against any proper local alternative.
Under the normality of errors, or more generally the conditions of Proposition 2, using that matrix $\frac{1}{\sqrt{\bar{q}}}Z^{*}$ is in the GOE for dimension TkT-k, i.e., $vech(Z^{*})\sim N(0,\bar{q}I_{p})$, we have $LR(k)\Rightarrow\bar{q}\chi^{2}(df,\lambda^{2}/\bar{q})$ from (4). The local power is a function solely of the squared Euclidean norm of the vector $vech(\Delta)$ measuring local distance from the $k$-factor specification, divided by $\bar{q}$.
@@ 6.2 AUMPI tests
[…]
@@
Matrices $\hat{G}$ and $\hat{\boldsymbol{D}}$ are both consistent up to post-multiplication by an orthogonal matrix. This point yields a group of orthogonal transformations under which we require the test statistics to be invariant, because the choices for the normalization of $\hat{G}$ and $\hat{\boldsymbol{D}}$ should be immaterial for the testing outcome. Here, we do not deal with invariance to data transformations but rather with invariance to parameterization of $\hat{G}$ and $\hat{\boldsymbol{D}}$. However, if we consider tests based on the elements of vector $\hat{W}$, this difference is immaterial. In Appendix B.6, we show that the maximal invariant under this group is provided by $\hat{W}^{\prime}\hat{W}=nvech(\hat{S}^{*})^{\prime}M_{\hat{\boldsymbol{X}}}vech(\hat{S}^{*})$. Since $\sqrt{n}vech(\hat{S}^{*})$ belongs to the range of matrix $M_{\hat{\boldsymbol{X}}}$, we have $\hat{W}^{\prime}\hat{W}=\frac{n}{2}\|\hat{S}^{*}\|^{2}=\frac{n}{2}\|\hat{S}\|^{2}$. Therefore, the invariant tests are functions of the squared norm of $\hat{S}$, which is asymptotically equivalent to the LR statistic (up to the factor $1/2$).
@@
In the Gaussian case, or more generally under the conditions of Proposition 2, the LR statistic follows asymptotically a scaled non-central chi-square distribution with $df$ degrees of freedom and non-centrality parameter $\lambda^{2}/\bar{q}=\sum_{j=1}^{df}\lambda_{j}^{2}$ as shown in the previous subsection. Thus, we can simplify the null and alternative hypotheses of our testing problem asymptotically and locally to a one-sided test with null hypothesis $H_{0}(k):\lambda^{2}=0$ vs. alternative hypothesis $H_{1,loc}(k):\lambda^{2}>0$. The scaling constant $q>0$ plays no role in the power analysis. It means that the LR test is an AUMPI test (Lehmann and Romano (2005) Chapters 3 and 13). Indeed, the density $g(z;df,\lambda^{2})$ of the $\chi^{2}(df,\lambda^{2})$ distribution is Totally Positive of order 2 ($TP2$) in $z$ and $\lambda^{2}$ (Eaton (1987) Example A.1 p. 468); see Miravete (2011) for a review of applications of TP2 in economics. A density, which is $TP2$ in $z$ and $\lambda^{2}$, has the Monotone Likelihood Ratio (MLR) property (Eaton (1987) p. 467). Since $g(z;df,\lambda^{2})/g(z;df,0)$ is an increasing function in $z$, it gives the AUMPI property. Bollerslev et al. (2025) also achieve to show the MLR property (with technical arguments markedly dfferent from the techniques commonly employed in high-frequency financial econometrics), and thus the AUMPI property for their Marubozu test.
@@
In the general case with $df>1$, when neither Gaussianity nor the conditions of Proposition 2 apply, we cannot use the same reasoning, since the density $f(z;\lambda_{1},...,\lambda_{df})$ of $\sum_{j=1}^{df}\mu_{j}\chi^{2}(1,\lambda_{j}^{2})$, with $\mu_{j}>0$, j=1,,dfj=1,...,df, is not a function of $\lambda^{2}=\sum_{j=1}^{df}\mu_{j}\lambda_{j}^{2}$ only, and thus cannot be $TP2$ in $z$ and $\lambda^{2}$. Instead, our strategy to get a new AUMPI result uses a power series representation of the density of $\sum_{j=1}^{df}\mu_{j}\chi^{2}(1,\lambda_{j}^{2})$ in terms of central chi-square densities from Kotz, Johnson, and Boyd (1967). Under the sufficient condition (5) in Proposition 4, the density ratio $\frac{f(z;\lambda_{1},...,\lambda_{df})}{f(z;0,...,0)}$ is monotone increasing in $z$.
@@
Let Assumptions 1-8 hold. (a) Let us assume that, for any DGP in the subset $\bar{H}_{1,loc}(k)\subset H_{1,loc}(k)$ of the local alternative hypothesis, we have for any integer $m\geq 3$: